# Decoding Chess Puzzle Play and Standard Cognitive Tasks for BCI: A Low-Cost EEG Study

### v1

@@ 4.3 Statistical Analysis of Cross-Task EEG Signal Differentiation
Table 8: ANOVA results from modeling EEG waveband data as a function of task type. To avoid multicollinearity between frequency bands, separate models were created for each waveband. P-values were adjusted using the Benjamini-Hochberg procedure to control for multiple comparisons. Significant differences in neural activity among tasks were observed across all frequency bands analyzed.
Band NumDF DenDF F value Pr($>$F) p.adj p.adj.sig $\eta_{p}^{2}$ 95% CI
$\theta$ 3.00 101.52 18.35 $<$0.001 $<$0.001 *** 0.35 [0.22,1]
$\alpha_{1}$ 3.00 100.10 24.48 $<$0.001 $<$0.001 *** 0.42 [0.30,1]
$\alpha_{2}$ 3.00 98.74 18.82 $<$0.001 $<$0.001 *** 0.36 [0.23,1]
$\beta_{1}$ 3.00 97.21 9.25 $<$0.001 $<$0.001 *** 0.22 [0.10,1]
$\beta_{2}$ 3.00 97.09 6.88 $<$0.001 $<$0.001 *** 0.18 [0.06,1]
$\gamma_{1}$ 3.00 96.95 8.56 $<$0.001 $<$0.001 *** 0.21 [0.09,1]
$\gamma_{2}$ 3.00 97.20 6.34 0.001 0.001 *** 0.16 [0.05,1]
Table 9: Post-hoc contrast results from modeling EEG waveband data as a function of task. Significant differences were discovered across a variety of frequency bands for varying tasks, indicating that differing neural signatures associated with the tasks are measurable with the MUSE 2 device. Among wavebands, effect sizes in the relatively low bands ($\theta$, $\alpha_{1}$, $\alpha_{2}$) are the largest, with a notable exceptions of increased power of Chess as compared to Stroop across all bands.
Band Task1 Task2 Est. SE df t p p.adj p.adj.sig $\eta_{p}^{2}$ 95% CI
$\theta$ Chess N-Back 0.18 0.04 100.35 4.91 $<$0.001 $<$0.001 *** 0.19 [0.09,1]
$\theta$ Chess Rotation -0.05 0.04 98.11 -1.37 0.524 0.667 ns 0.02 [0.00,1]
$\theta$ Chess Stroop 0.18 0.04 112.30 4.40 $<$0.001 $<$0.001 *** 0.15 [0.06,1]
$\theta$ N-Back Rotation -0.23 0.04 98.07 -6.19 $<$0.001 $<$0.001 *** 0.28 [0.16,1]
$\theta$ N-Back Stroop -0.00 0.04 $1$0.12 0.999 0.999 ns 0.00 [0.00,1]
$\theta$ Rotation Stroop 0.23 0.04 109.75 5.72 $<$0.001 $<$0.001 *** 0.23 [0.12,1]
$\alpha_{1}$ Chess N-Back 0.16 0.03 99.32 5.76 $<$0.001 $<$0.001 *** 0.25 [0.14,1]
$\alpha_{1}$ Chess Rotation -0.03 0.03 97.18 -1.11 0.682 0.842 ns 0.01 [0.00,1]
$\alpha_{1}$ Chess Stroop 0.16 0.03 110.31 5.51 $<$0.001 $<$0.001 *** 0.22 [0.11,1]
$\alpha_{1}$ N-Back Rotation -0.19 0.03 98.16 -6.78 $<$0.001 $<$0.001 *** 0.32 [0.20,1]
$\alpha_{1}$ N-Back Stroop 0.01 0.03 111.33 0.20 0.997 0.999 ns 0.00 [0.00,1]
$\alpha_{1}$ Rotation Stroop 0.19 0.03 109.25 6.61 $<$0.001 $<$0.001 *** 0.29 [0.17,1]
$\alpha_{2}$ Chess Stroop 0.16 0.03 110.38 5.95 $<$0.001 $<$0.001 *** 0.24 [0.14,1]
$\alpha_{2}$ Chess N-Back 0.13 0.02 99.30 5.37 $<$0.001 $<$0.001 *** 0.23 [0.12,1]
$\alpha_{2}$ Chess Rotation 0.01 0.03 97.14 0.39 0.979 0.999 ns 0.00 [0.00,1]
$\alpha_{2}$ N-Back Rotation -0.12 0.03 98.13 -4.88 $<$0.001 $<$0.001 *** 0.20 [0.09,1]
$\alpha_{2}$ N-Back Stroop 0.03 0.03 111.40 1.01 0.743 0.891 ns 0.01 [0.00,1]
$\alpha_{2}$ Rotation Stroop 0.15 0.03 109.32 5.63 $<$0.001 $<$0.001 *** 0.22 [0.12,1]
$\beta_{1}$ Chess N-Back 0.12 0.04 101.32 2.90 0.023 0.049 * 0.08 [0.01,1]
$\beta_{1}$ Chess Rotation 0.08 0.04 99.96 1.94 0.217 0.304 ns 0.04 [0.00,1]
$\beta_{1}$ Chess Stroop 0.23 0.04 107.61 5.41 $<$0.001 $<$0.001 *** 0.21 [0.11,1]
$\beta_{1}$ N-Back Rotation -0.04 0.04 99.80 -0.89 0.809 0.928 ns 0.01 [0.00,1]
$\beta_{1}$ N-Back Stroop 0.12 0.04 107.44 2.73 0.037 0.070 ns 0.06 [0.01,1]
$\beta_{1}$ Rotation Stroop 0.15 0.04 105.98 3.58 0.003 0.007 ** 0.11 [0.03,1]
$\beta_{2}$ Chess N-Back 0.05 0.06 102.33 0.88 0.817 0.928 ns 0.01 [0.00,1]
$\beta_{2}$ Chess Rotation 0.16 0.06 101.38 2.70 0.040 0.073 ns 0.07 [0.01,1]
$\beta_{2}$ Chess Stroop 0.26 0.06 106.56 4.32 $<$0.001 0.001 ** 0.15 [0.06,1]
$\beta_{2}$ N-Back Rotation 0.11 0.06 100.58 1.86 0.252 0.342 ns 0.03 [0.00,1]
$\beta_{2}$ N-Back Stroop 0.21 0.06 105.76 3.53 0.003 0.008 ** 0.11 [0.03,1]
$\beta_{2}$ Rotation Stroop 0.11 0.06 104.62 1.74 0.308 0.404 ns 0.03 [0.00,1]
$\gamma_{1}$ Chess N-Back 0.13 0.05 102.04 2.47 0.071 0.111 ns 0.06 [0.01,1]
$\gamma_{1}$ Chess Rotation 0.14 0.05 100.95 2.62 0.049 0.083 ns 0.06 [0.01,1]
$\gamma_{1}$ Chess Stroop 0.30 0.06 106.91 5.23 $<$0.001 $<$0.001 *** 0.20 [0.10,1]
$\gamma_{1}$ N-Back Rotation 0.01 0.05 100.34 0.21 0.997 0.999 ns 0.00 [0.00,1]
$\gamma_{1}$ N-Back Stroop 0.17 0.06 106.28 2.95 0.020 0.044 * 0.08 [0.01,1]
$\gamma_{1}$ Rotation Stroop 0.15 0.06 105.03 2.75 0.035 0.070 ns 0.07 [0.01,1]
$\gamma_{2}$ Chess N-Back 0.13 0.06 102.21 2.20 0.130 0.195 ns 0.05 [0.00,1]
$\gamma_{2}$ Chess Rotation 0.12 0.06 101.12 1.97 0.205 0.297 ns 0.04 [0.00,1]
$\gamma_{2}$ Chess Stroop 0.28 0.06 107.11 4.52 $<$0.001 $<$0.001 *** 0.16 [0.07,1]
$\gamma_{2}$ N-Back Rotation -0.01 0.06 100.35 -0.17 0.998 0.999 ns 0.00 [0.00,1]
$\gamma_{2}$ N-Back Stroop 0.15 0.06 106.33 2.48 0.069 0.111 ns 0.05 [0.01,1]
$\gamma_{2}$ Rotation Stroop 0.16 0.06 105.05 2.65 0.046 0.080 ns 0.06 [0.01,1]
@@
From these data a few notable trends arise. Firstly are the similarities within the power disparities in the low-to-middle frequency bands ($\theta$, $\alpha_{1}$, and $\alpha_{2}$) between two groups of two tasks: Chess and Rotation, both significantly higher in all such bands (see Table 9) than the N-Back and Stroop tasks. This clustering suggests these tasks may share underlying cognitive mechanisms, particularly in how they engage spatial processing and mental manipulation of visual information.
@@
Considering this more deeply, as compared to Stroop and N-Back, both Chess and mental rotation heavily recruit visuospatial working memory systems. Although increases in $\alpha$ power have been found in relation to event-related desynchronization (ERS), other work has demonstrated that under certain conditions $\alpha$ power increases during task activation (Klimesch et al., 2007), specifically in a visuospatial manipulation task as compared to a memory task Sauseng et al. (2005). Further, frontal $\alpha$ power has been linked to creative ideation reflecting internal processing Fink and Benedek (2014), which largely would be exhibited during the Chess task.
@@
Elevated $\theta$ power likely reflects an increase of overall mental effort Smit et al. (2005) and on executive control processes Cavanagh and Frank (2014); Sauseng et al. (2010) required by these tasks during the manipulation of complex spatial representations. Chess positions and mental rotation problems both require the generation and manipulation of spatial configurations within working memory, processes known to engage frontoparietal networks that communicate through synchronized $\theta$ oscillations Sauseng et al. (2010). In the case of Chess puzzles, long-term memory structures will be engaged in the attempt to recall puzzle patterns previously encountered; in the case of mental rotation, as the visual attention is limited to a single block at a time, there is an implicit need to store and remember the block not currently attended to. Although the increase in $\theta$ power during increased levels of the N-Back task indicate increased storage and recall of information, the relative requirement of overall level-independent (i.e. all levels of N-Back) storage and recall is less than the additional neural processes required by Chess and Rotation, including manipulation of relatively larger quantities of visual information, and long-term informational interfacing in the case of Chess puzzles Sauseng et al. (2010). Indeed, N-Back and Stroop tasks involve rapid (1s) sequential trials, which can provide targeted tasks to explore specific neurophysiological dynamics, but overall may involve less cognitive power than the activity of the Chess and Rotation tasks.
@@
This distinction also maps to Baddeley’s multicomponent model of working memory Baddeley (2000): Chess and Rotation more heavily engage the visuospatial sketchpad component, whereas N-Back and Stroop by contrast more heavily engage the phonological loop.
@@
The pattern between Chess, Rotation, and Stroop persists in the $\beta_{1}$ band, with a notable exception: the N-Back task is not significantly lower than either Chess or Rotation. This divergence likely stems from the N-Back task’s intensive working memory demands, which have been shown to relate to bursts of activity in $\beta$ band power Lundqvist et al. (2018).
@@
Moving into the higher frequency bands, N-Back and Chess show higher power than Stroop in the $\beta_{2}$ band. This result may potentially be a consequence of the decreased $\beta_{2}$ activation due to conflict inhibition discussed in the within-task results for Stroop, combined with increased mental workload exhibited by N-Back. Chess continues to have significantly higher power than Stroop in the highest bands ($\beta_{2}$, $\gamma_{1}$, and $\gamma_{2}$). This comparison is likely due to the engagement of integrated and complex cognitive processes in the Chess task Atherton et al. (2003). Yet, although Chess represents the consistently highest total amount of $\beta$ and $\gamma$, no significant differences among Chess, N-Back, and Rotation were discovered in these bands, indicating that identifying the particular nuances of the differential nature of the complexity of the Chess task is more challenging. Inspection of distinctions among more areas of the brain may yield insights into the full complexities of Chess as compared to Rotation and N-Back.
@@
4.4 Within-Task Workload Machine Learning Results
As shown in Table 10, the within-task workload classification presented significant challenges. The N-Back task achieved a 63% F1-score, suggesting some promise, while other tasks performed near chance levels when using the MUSE 2 device for online classification. Several factors may explain these results. First, our limited dataset size may have constrained the models’ learning capabilities.
Secondly, the N-Back task inherently induces the most distinct cognitive load differences between difficulty levels Brouwer et al. (2012). By contrast, although the Stroop task can reliably induce cognitive interference measurable at the macro level, the subtleties of the neural signatures elicited between the congruent and incongruent states may be difficult to capture on a per-trial basis with the resolution of the MUSE 2 device Krigolson et al. (2017). Likewise, although larger degrees of mental rotation take more time on average to complete, the neural signatures underlying the tasks may not be sufficiently distinct to measure with the MUSE 2. In the case of both Stroop and rotation, it is also possible that complex networks extending beyond the prefrontal region (in particular, parietal and visuospatial networks) may be required to interpret the data more clearly. Regarding Chess puzzles, it is possible that at the higher degrees of difficulty, time pressure may have introduced artifacts into the neural signals which might have affected classification performance.
@@ 4.4 Within-Task Workload Machine Learning Results
Table 10: Task-specific classification of cognitive workload using machine learning. N-Back performs the best, with 63% F1-score over 2-class classification. Chess performs just better than chance, with 54%. The Rotation and Stroop task classifiers do not perform better than chance. Note that the results are averages over all Monte-Carlo iterations.
@@
4.5 Cross-Task Machine Learning Results
Full results are shown in Table 11. The machine learning model for between-task classification demonstrated robust performance, achieving a macro-average F1-score of 49% (95% confidence interval of [0.493, 0.495]) compared to the expected random chance performance of 25%. The Rotation task showed the strongest classification metric with an F1-score of 53%, while Chess, Stroop, and N-Back tasks achieved F1-scores of 50%, 49%, and 45%, respectively. This performance is particularly noteworthy given the inherent challenges of EEG classification. The balanced precision and recall scores across all tasks (ranging from approximately 46% to 56%) indicate consistent and reliable classification capabilities. The relatively similar performance levels of the tasks align with our earlier findings regarding unique combinations of neural activation patterns discussed earlier. These classification results demonstrate that the MUSE 2 device can effectively differentiate between distinct cognitive tasks, with performance approximately twice that of random chance across all task types.
Table 11: Cross-task classification of EEG signals using machine learning. These results are averages across all Monte Carlo iterations.
@@ 4.5 Cross-Task Machine Learning Results
5 Future Work and Methodological Considerations
5.1 Signal Processing Enhancements
@@
5.2 Advanced Machine Learning Applications
Recent developments in deep learning architectures suggest promising directions for improving workload classification. (Craik et al., 2019) demonstrates deep learning models’ success in handling temporal dynamics inherent in EEG data, while (Hossain et al., 2023) shows transformer-based architectures can effectively capture long-range dependencies in EEG signals. Building on methodological insights from (Murungi et al., 2023), future work could implement attention mechanisms and transfer learning to address our dataset limitations, particularly beneficial for complex tasks like Chess where traditional frequency-domain features may not fully capture neural dynamics. Future investigations should balance architectural sophistication with practical implementation constraints for consumer-grade EEG devices. This aligns with current trends toward deployable BCI systems that maintain both accuracy and computational efficiency (Hossain et al., 2023). Additionally, incorporating self-supervised learning techniques could address data limitation challenges (Craik et al., 2019), potentially improving classification robustness across cognitive tasks through more effective utilization of unlabeled EEG data.
5.3 Sample Size and Generalizability
@@
5.4 Task Difficulty Calibration
@@ 6 Probe Limitations
7 Signal Connectivity Issues
@@
8 Timing Limitations for the Chess Task
The time limit posed in the Chess task of 30 seconds per puzzle enabled collection of a maximal number of puzzles, however the factor of time pressure may also have disrupted natural cognitive processes. We recommend future work which compares neural responses under both timed and untimed conditions during similar Chess puzzle scenarios.
9 Conclusion
@@
The research code used to run the experimental tasks can be accessed at https://github.com/mattrussell2/chess-mw-MUSE. The preprocessed EEG datasets can be found at https://github.com/mattrussell2/chess-mw-MUSE-DATA. The chess puzzles database is not maintained by the study team, and can instead be downloaded from https://database.lichess.org/ (Lichess Team, 2024a).
Conflict of Interest Statement
We declare that this research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.
Author Contributions
MR: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Project administration, Software, Supervision, Validation, Visualization, Writing - original draft, Writing - review & editing; SY: Conceptualization, Investigation, Methodology, Writing - original draft, Writing - review & editing; WX: Conceptualization, Investigation, Methodology, Software, Writing - review & editing; KZ: Conceptualization, Investigation, Methodology, Project administration, Writing - original draft, Writing - review & editing; AS: Conceptualization, Investigation, Writing - original draft, Writing - review & editing; RJ: Conceptualization, Funding acquisition, Project administration, Supervision, Writing - review & editing.
Funding
Google, Inc.
Acknowledgments
Many thanks to Soraya Basrai, Hannah Jiang, Edward Lai, Ryan Poon, Rohan Dugad, and Iris Yang for their help along the way.

### v2

@@ IV-C Statistical Analysis of Cross-Task EEG Signal Differentiation
From these data a few notable trends arise. Firstly are the similarities within the power disparities in the low-to-middle frequency bands ($\theta$, $\alpha_{1}$, and $\alpha_{2}$) between two groups of two tasks: Chess and Rotation, both significantly higher in all such bands (see LABEL:supp-tab:BETWEEN_TASK_PHOCS in the supplementary material) than the N-Back and Stroop tasks. This clustering suggests these tasks may share underlying cognitive mechanisms, particularly in how they engage spatial processing and mental manipulation of visual information.
@@
Considering this more deeply, as compared to Stroop and N-Back, both Chess and mental rotation heavily recruit visuospatial working memory systems. Although increases in $\alpha$ power have been found in relation to event-related desynchronization (ERS), other work has demonstrated that under certain conditions $\alpha$ power increases during task activation [69], specifically in a visuospatial manipulation task as compared to a memory task [70]. Further, frontal $\alpha$ power has been linked to creative ideation reflecting internal processing [71], which largely would be exhibited during the Chess task.
@@
Elevated $\theta$ power likely reflects an increase of overall mental effort [72] and on executive control processes [73, 74] required by these tasks during the manipulation of complex spatial representations. Chess positions and mental rotation problems both require the generation and manipulation of spatial configurations within working memory, processes known to engage frontoparietal networks that communicate through synchronized $\theta$ oscillations [74]. In the case of Chess puzzles, long-term memory structures will be engaged in the attempt to recall puzzle patterns previously encountered; in the case of mental rotation, as the visual attention is limited to a single block at a time, there is an implicit need to store and remember the block not currently attended to. Although the increase in $\theta$ power during increased levels of the N-Back task indicate increased storage and recall of information, the relative requirement of overall level-independent (i.e. all levels of N-Back) storage and recall is less than the additional neural processes required by Chess and Rotation, including manipulation of relatively larger quantities of visual information, and long-term informational interfacing in the case of Chess puzzles [74]. Indeed, N-Back and Stroop tasks involve rapid (1s) sequential trials, which can provide targeted tasks to explore specific neurophysiological dynamics, but overall may involve less cognitive power than the activity of the Chess and Rotation tasks.
@@
This distinction also maps to Baddeley’s multicomponent model of working memory [75]: Chess and Rotation more heavily engage the visuospatial sketchpad component, whereas N-Back and Stroop by contrast more heavily engage the phonological loop.
@@
The pattern between Chess, Rotation, and Stroop persists in the $\beta_{1}$ band, with a notable exception: the N-Back task is not significantly lower than either Chess or Rotation. This divergence likely stems from the N-Back task’s intensive working memory demands, which have been shown to relate to bursts of activity in $\beta$ band power [76].
@@
Moving into the higher frequency bands, N-Back and Chess show higher power than Stroop in the $\beta_{2}$ band. This result may potentially be a consequence of the decreased $\beta_{2}$ activation due to conflict inhibition discussed in the within-task results for Stroop, combined with increased mental workload exhibited by N-Back. Chess continues to have significantly higher power than Stroop in the highest bands ($\beta_{2}$, $\gamma_{1}$, and $\gamma_{2}$). This comparison is likely due to the engagement of integrated and complex cognitive processes in the Chess task [77]. Yet, although Chess represents the consistently highest total amount of $\beta$ and $\gamma$, no significant differences among Chess, N-Back, and Rotation were discovered in these bands, indicating that identifying the particular nuances of the differential nature of the complexity of the Chess task is more challenging. Inspection of distinctions among more areas of the brain may yield insights into the full complexities of Chess as compared to Rotation and N-Back.
@@
IV-D Within-Task Workload Machine Learning Results
As shown in Table I, the within-task workload classification presented significant challenges. The N-Back task achieved a 63% F1-score, suggesting some promise, while other tasks performed near chance levels when using the MUSE 2 device for online classification. Several factors may explain these results. First, our limited dataset size may have constrained the models’ learning capabilities.
Secondly, the N-Back task inherently induces the most distinct cognitive load differences between difficulty levels [64]. By contrast, although the Stroop task can reliably induce cognitive interference measurable at the macro level, the subtleties of the neural signatures elicited between the congruent and incongruent states may be difficult to capture on a per-trial basis with the resolution of the MUSE 2 device [16]. Likewise, although larger degrees of mental rotation take more time on average to complete, the neural signatures underlying the tasks may not be sufficiently distinct to measure with the MUSE 2. In the case of both Stroop and rotation, it is also possible that complex networks extending beyond the prefrontal region (in particular, parietal and visuospatial networks) may be required to interpret the data more clearly. Regarding Chess puzzles, it is possible that at the higher degrees of difficulty, time pressure may have introduced artifacts into the neural signals which might have affected classification performance.
@@ IV-D Within-Task Workload Machine Learning Results
TABLE I: Task-specific classification of cognitive workload using machine learning. N-Back performs the best, with 63% F1-score over 2-class classification. Chess performs just better than chance, with 54%. The Rotation and Stroop task classifiers do not perform better than chance. Note that the results are averages over all Monte-Carlo iterations.
@@
IV-E Cross-Task Machine Learning Results
Full results are shown in Table II. The machine learning model for between-task classification demonstrated robust performance, achieving a macro-average F1-score of 49% (95% confidence interval of [0.493, 0.495]) compared to the expected random chance performance of 25%. The Rotation task showed the strongest classification metric with an F1-score of 53%, while Chess, Stroop, and N-Back tasks achieved F1-scores of 50%, 49%, and 45%, respectively. This performance is particularly noteworthy given the inherent challenges of EEG classification. The balanced precision and recall scores across all tasks (ranging from approximately 46% to 56%) indicate consistent and reliable classification capabilities. The relatively similar performance levels of the tasks align with our earlier findings regarding unique combinations of neural activation patterns discussed earlier. These classification results demonstrate that the MUSE 2 device can effectively differentiate between distinct cognitive tasks, with performance approximately twice that of random chance across all task types.
TABLE II: 4-way cross-task classification (random chance level is 25%) of EEG signals using machine learning. These results are averages across all Monte Carlo iterations.
@@ IV-E Cross-Task Machine Learning Results
V Future Work and Methodological Considerations
V-A Signal Processing Enhancements
@@
V-B Advanced Machine Learning Applications
Recent developments in deep learning architectures suggest promising directions for improving workload classification. [78] demonstrates deep learning models’ success in handling temporal dynamics inherent in EEG data, while [79] shows transformer-based architectures can effectively capture long-range dependencies in EEG signals. Building on methodological insights from [80], future work could implement attention mechanisms and transfer learning to address our dataset limitations, particularly beneficial for complex tasks like Chess where traditional frequency-domain features may not fully capture neural dynamics. Future investigations should balance architectural sophistication with practical implementation constraints for consumer-grade EEG devices. This aligns with current trends toward deployable BCI systems that maintain both accuracy and computational efficiency [79]. Additionally, incorporating self-supervised learning techniques could address data limitation challenges [78], potentially improving classification robustness across cognitive tasks through more effective utilization of unlabeled EEG data.
V-C Sample Size and Generalizability
@@
V-D Task Difficulty Calibration
@@ V-E Probe Limitations
V-F Signal Connectivity Issues
@@
VI Conclusion
@@
The research code used to run the experimental tasks can be accessed at REDACTED.
The preprocessed EEG datasets can be found at REDACTED. The chess puzzles database is not maintained by the study team, and can instead be downloaded from https://database.lichess.org/ [36].